# SynthesizRR: Generating Diverse Datasets with Retrieval Augmentation

### v1

@@ 4 Experimental Setup
Tasks and complexity. We experiment on 6 datasets described in Table 1, which are selected to give diverse task inversion steps. Previous work primarily benchmarks sentiment and topic classification datasets like AG News (Zhang et al., 2015), NYT (Meng et al., 2019), and Reddit (Geigle et al., 2021). We broaden from topic classification, which primarily involves summarization during the task inversion step, which LLMs are adept at (Goyal et al., 2022). Hyperpartisan (Kiesel et al., 2019) detects bias in political news, so the task inversion step includes a more substantial rewriting of neutral retrieved articles to form biased examples. Task inversion for Humor (Ziser et al., 2020), Category Blitzer et al. (2007) and Polarity (Ni et al., 2019) involves generating humorous questions and reviews from retrieved product details, which require additional skills.
Corpus Domain Docs Item Token
@@
Table 2: Corpus statistics. Used LLaMa2 tokenizer.
Each task in Table 1 has at least one corpus which can be considered in-domain for the task. We ground the synthesis to this in-domain corpus during SynthesizRR. Prompts for all tasks are in Appendix E.
In §7, we compare to six prior approaches on four popular tasks: Yelp Polarity (Zhang et al., 2015), IMDb Maas et al. (2011), SST-2 Socher et al. (2013) and AG News, when grounding to smaller corpora (42k-150k documents). We view these tasks as less aligned with the goals our paper, so we do not focus our evaluation on them here.
@@
Gold There is decent bass, but the highs are a bit soft. A quick tweak to my equalizer, and they’re great. After reading several of the reviews on Amazon, I was a bit worried about the sound, but now that I have them I’m very happy. They’re a good price, and sooooo much better than the little ipod-like earbuds I’ve tried before. Those never stayed in my ear, and the bass never made me happy
@@
Retrieval corpora. Table 2 describes corpora used for retrieval. We consider four corpora in different domains, each with millions of records: RealNews/Dominant (US/EU News), RealNews/Regional (Regional News), RealNews/India (Indian News), Products (Amazon products metadata (Ni et al., 2019)).
@@
Retrieval model. We use Contriever (Izacard et al., 2022) for dense retrieval from each corpus. This performs a semantic match between the query and each document using cosine-similarity. In Appendix B, we also perform an ablation study using sparse retriever BM25, which does lexical matching between each query-document pair.
@@
Teacher LLMs. As teacher models, we primarily use a frozen LLaMa-2 Chat 13B (Touvron et al., 2023) for the task inversion step in SynthesizRR and FewGen. We also experiment with ClaudeV1 (Appendix H). For in-context learning (ICL) Brown et al. (2020), we select examples randomly from the train set: 50 ICL examples/class for multi-class and 100/class for binary tasks. We believe this is a realistic number of examples that a system designer could source if they were to put some effort into building a specialist model. We pick these numbers to provide sufficient ICL diversity in 32-shot FewGen and SynthesizRR.
@@
Student model. Specialization performance is measured on student LM DeBERTa-v3-Large (435M params, He et al. (2021)).
@@
Figure 3: Self-BLEU ()($\downarrow$)( ↓ ) for ngrams n=1-5. Comparison: Gold, FewGen 0-shot, FewGen 32-shot, SynthesizRR 0-shot, SynthesizRR 3-shot Retr-ICL, SynthesizRR 32-shot No-Retr-ICL.
@@
Evaluation criteria. Text generation can be challenging to evaluate objectively in multi-task scenarios (Chang et al., 2024). Therefore in §5 we evaluate synthetic text based on several criterion, to detect behaviours we observe during synthesis as in Table 3. Self-BLEU (Papineni et al., 2002; Zhu et al., 2018) measures lexical diversity of the dataset based on $n$-gram overlap between pairs of examples. Entity entropy measures the diversity of entities using the probability distribution of each of 16 entity-types (inferred from a pretrained named entity recognition model). Dataset which over-represent popular entities score lower on entropy. On the other hand, Entity recall and Entity KL divergence compares the similarity of entities compared to Gold, and datasets which reproduce entities frequently seen in Gold data score higher. MAUVE (Liu et al., 2021) measure similarity to human-written text using pretrained representations from a GPT2-XL model; this indicates distributional differences that would be perceived by language models.
Besides the final accuracy of the student model, we also consider label preservation accuracy, which is obtained from an “oracle” model for the task. We construct the oracle from Gold data by running a grid-search over DeBERTa-v3-Large hyperparams, splitting 80% of the Gold train set for fine-tuning and 20% for validation. Then, we measure the fraction of synthetic examples which the oracle classifies to belong to the prompted target class. This indicates the adherence of the generated example to the class it should belong to, as per the prompt.
@@
Figure 4: Entity entropy ()($\uparrow$)( ↑ ) on ToI (headlines) and Category (reviews). Comparison: Gold, FewGen 32-shot, SynthesizRR 3-shot Retr-ICL and SynthesizRR 32-shot No-Retr-ICL. Zero-shot results are similar for SynthesizRR and worse for FewGen; we omit them.
5 Results: Intrinsic Evaluation
In this section we focus on evaluating intrinsic properties of the generated datasets, including their diversity and entity coverage. We focus on a LLaMa-2 Chat 13B teacher LLM. We generate datasets of a fixed size comparable to the number of Gold rows: $8$k rows (AG News, ToI Headlines, Category), $4$k rows (Polarity) or $2$k rows (Hyperpartisan (516 originally), Humor). Examples of generations on each dataset can be found in Appendix F.
RQ: Does SynthesizRR address entity diversity? Popularity bias is a phenomenon wherein LLM generations tend to over-represent popular “head” entities. It has been studied for QA tasks (Mallen et al., 2023; Kandpal et al., 2023).
In Figure 4 we see how SynthesizRR eliminates popularity bias in several entity types at once. Here, we use Spacy NER model en_core_web_lg (Honnibal et al., 2020) to detect entity occurrences in each synthetic example. By sourcing from the long-tail of retrieval results ($k=50$), the generated dataset has much higher entity entropy, meaning the probability distribution of each entity-type is closer to a uniform distribution compared to FewGen. This positions SynthesizRR closer to Gold, which also shows high entity entropy.
@@ 5 Results: Intrinsic Evaluation
SynthzRR 319 3991 1989 397 0.35 2.35
SynthzRR 314 2699 1464 363 0.32 2.52
Table 4: Entity similarity in Category ($8$K). We show the counts of unique and total entities for 4 entity-types. Entity recall measures the fraction of Gold entities co-occuring in the synthetic data; in the bottom half, we additionally weigh each entity by its frequency in Gold. Notation: *32-shot; 3-shot Retr-ICL; 32-shot No-Retr-ICL.
RQ: How is entity similarity in synthetic data affected by grounding to an in-domain corpus? Generating diverse entities is not beneficial if they are not also present in Gold. We thus study entity similarity. For the Category task we generate $8$k product reviews and randomly select $8$k Gold examples. In Table 4, we measure entity recall, and find that the occurrence of Gold entities is 100%-140% higher in SynthesizRR than FewGen. The KL divergence of each entity-distribution is also lower. We finally consider the entity coverage (unique entities) and entity density (total entities). Compared to Gold, FewGen tends to produce fewer unique entities (places, events, languages, currencies, etc). Each FewGen example also has a lower density of entities, as visible in Table 3. SynthesizRR coverage and density more closely match Gold.
RQ: How distributionally similar are our generated examples and human-written examples? We calculate MAUVE score, which uses GPT-2XL (Radford et al., 2019) to embed texts and compare them based on Type-1 and Type-2 errors (Liu et al., 2021). We see from Table 5 that zero-shot generations are quite dissimilar in both approaches compared to few-shot methods.
@@
SynthzRR 92.0 72.8 87.9 75.2 87.5 89.9
SynthzRR $91.76$ $67.86$ $67.16$ $75.14$ $86.95$ $83.2$
Table 5: MAUVE similarity score ($\uparrow$) using GPT2-XL embeddings. Notation: *32-shot; 3-shot Retr-ICL; 32-shot No-Retr-ICL.
@@
(Dataset size) ($8$K) ($4$K) ($8$K) ($8$K) ($2$K) ($4$K)
Gold 91.015 78991.01578991.015\,78991.015 789 93.230 76993.23076993.230\,76993.230 769 $82.488$ 81.516 66781.51666781.516\,66781.516 667 $93.06$ $95.2645$ 89.43
Zero shot
@@
Table 6: Test Accuracy ($\uparrow$) after distilling DeBERTa-v3-Large student from LLaMa-2 Chat 13B and Claude Instant-v1. We report mean of 5 runs and rerun all 5 in cases of std $\geq$ 6.0. The top half considers zero-shot synthesis and bottom half uses in-context learning, where we bold the best result under each paradigm. Notation: *32-shot; 3-shot Retr-ICL; 32-shot No-Retr-ICL.
Surprisingly, SynthesizRR generations are much more similar to human text than FewGen. This is surprising as nothing in our content sourcing strategy explicitly guides SynthesizRR generations to match the distribution of Gold. We thus manually inspect generations and discover an interesting pattern which can be attributed to content sourcing. As shown earlier, and in Table 3, the density of entities is higher under SynthesizRR. FewGen produces generations which obey the prompt, but are very bland and do not include specifics. On the other hand, by obtaining information-rich documents, SynthesizRR is able to ground the task inversion step in details of the retrieved article/product. We hypothesise that this improves the MAUVE score towards Gold, which is similarly grounded in specifics.
We further analyze the effect of the retrieval corpus in Appendix C.
@@ 6 Results: Student distillation
Table 6 shows the results of training a DeBERTa-v3L student on generated datasets across several settings. We report the mean of 5 runs and rerun if the standard deviation is $\geq$6(indicating one or more models failed to train). In the zero-shot setting, we find that SynthesizRR performs much better than FewGen, despite using the same frozen teacher LLM. Note that SynthesizRR uses in-context examples for retrieval here whereas FewGen does not; our method has some additional supervision here. However, in this setting, we see clear gains during the task inversion stage ($\uparrow$12for LLaMa2 and $\uparrow$17.6for ClaudeV1). Thus, having access to retrieval yields a better final dataset, almost on par with the few-shot generation.
@@
With in-context learning, 3-shot SynthesizRR using the No-Retr-ICL strategy trains better students than 32-shot FewGen ($\uparrow$1.3for LLaMa2 and $\uparrow$3.2for ClaudeV1). We conclude that naively adding ICL examples is not an effective use of the LLM’s context window. Instead, a better content sourcing strategy positively influences the student distillation, leading to faster synthesis and better test performance.

### v3

@@ 4 Experimental Setup
Tasks and their difficulty. We perform our main experiments on the first 6 datasets in Table 1, selected carefully to measure how the teacher LLM performs on task inversion tasks of varying difficulty. Previous work only benchmarked sentiment and topic classification datasets like IMDb Maas et al. (2011) and AG News (Zhang et al., 2015). We broaden from topic classification, which primarily involves summarization during the task inversion step, which LLMs are adept at (Goyal et al., 2022). Hyperpartisan (Kiesel et al., 2019) detects bias in political news, so the task inversion step includes a more substantial rewriting of neutral retrieved articles to form biased examples. Category and Polarity are prevalent product review tasks (Yu et al., 2023a, b; Gao et al., 2023); we generate reviews from retrieved products which must conform to categorical and sentiment classes. Task inversion for Humor (Ziser et al., 2020) involves generating humorous questions from retrieved product details, which requires additional skills from the teacher. Prompts for all tasks are in Appendix G.
Corpus Domain Size Doc. Tokens
@@
[…]
Table 2: Corpus statistics with LLaMa2 tokenizer.
Table 2 describes corpora used for retrieval. We consider five corpora in different domains, each with varying numbers of records. Three are subsets of RealNews (Zellers et al., 2019), as described in Appendix I: RealNews/Dominant (US/EU News), RealNews/Regional (Regional News), RealNews/India (Indian News). We also use Products (Amazon products metadata, (Ni et al., 2019)) and Movie Summary (movie summaries, (Bamman et al., 2013). Each task in Table 1 is associated with the corpus we consider most relevant. In §7, we compare to four prior approaches on three other tasks: IMDb Maas et al. (2011), SST-2 Socher et al. (2013) and AG News. These sentiment and topic tasks are less aligned with our goals and thus excluded from our main evaluation.
@@
Gold There is decent bass, but the highs are a bit soft. A quick tweak to my equalizer, and they’re great. After reading several of the reviews on Amazon, I was a bit worried about the sound, but now that I have them I’m very happy. They’re a good price, and sooooo much better than the little ipod-like earbuds I’ve tried before. Those never stayed in my ear, and the bass never made me happy.
@@
Models. We use Contriever (Izacard et al., 2022) for dense retrieval from each corpus. This performs a semantic match between the query and each document using cosine-similarity. In Appendix E, we also perform an ablation study using BM25 as a sparse retriever, which does lexical matching between each query-document pair.
@@
As teacher models, we primarily use a frozen Llama-2 Chat 13B (Touvron et al., 2023b) for the task inversion step in SynthesizRR and FewGen. We also experiment with Claude Instant-v1 as described in Appendix J. For in-context learning (ICL) Brown et al. (2020), we select examples randomly from the train set: 50 ICL examples/class for multi-class and 100/class for binary tasks. We believe this is a realistic number of examples that a system designer could source if they were to put some effort into building a specialist model. We explore approaches to bootstrap this seed set in limited-supervision settings Appendix C.
@@
Specialization performance is measured on student LMs DeBERTa-v3-Large (435M params, He et al. (2021)) and DistilBERT (66M params, Sanh et al. (2019)).
@@
Figure 3: Self-BLEU ()($\downarrow$)( ↓ ) for ngrams n=1-5. Comparison: Gold, FewGen 0-shot, FewGen 32-shot, SynthesizRR 0-shot, SynthesizRR 3-shot RetrICL, SynthesizRR 32-shot Non-RetrICL.
@@
Evaluation criteria. Text generation can be challenging to evaluate objectively in multi-task scenarios (Chang et al., 2024). Therefore in §5 we evaluate synthetic text based on several criterion, to detect behaviours we observe during synthesis as in Table 3. Self-BLEU (Papineni et al., 2002; Zhu et al., 2018) measures lexical diversity of the dataset based on $n$-gram overlap between pairs of examples. Entity entropy measures the diversity of entities using the probability distribution of each of 16 entity-types, inferred using spaCy’s en_core_web_lg (Honnibal et al., 2020). Datasets which over-represent popular entities score lower on entropy. On the other hand, Entity recall and Entity KL divergence compares the similarity of entities compared to Gold, and datasets which reproduce entities frequently seen in Gold data score higher. MAUVE (Liu et al., 2021) measures similarity to human-written text by using pretrained representations from a gpt2-xl model, indicating distributional differences in the generated text.
@@
Figure 4: Entity entropy ()($\uparrow$)( ↑ ) on ToI (headlines) and Category (reviews). Comparison: Gold, FewGen 32-shot, SynthesizRR 3-shot RetrICL and SynthesizRR 32-shot Non-RetrICL. Zero-shot results are similar for SynthesizRR and worse for FewGen; we omit them.
@@
SynzthRR 319 3991 1989 397 0.35 2.35
SynzthRR 314 2699 1464 363 0.32 2.52
Table 4: Entity similarity in Category ($8$K). We show the counts of unique and total entities for 4 entity-types. Entity recall measures the fraction of Gold entities co-occuring in the synthetic data; in the bottom half, we additionally weigh each entity by its frequency in Gold. Notation: *32-shot; 3-shot RetrICL; 32-shot Non-RetrICL.
5 Results: Intrinsic Evaluation
In this section, we focus on evaluating intrinsic properties of the generated datasets, including their diversity and entity coverage. We focus on a LLaMa-2 Chat 13B teacher LLM, retrieving from Contriever using corpora per Table 1 (we analyze changing the retrieval corpus in  Appendix D). We generate datasets of size in relation to the number of Gold rows: $8$K rows (AG News, ToI Headlines, Category), $4$K rows (Polarity) or $2$K rows (Hyperpartisan, Humor). Example generations are in Appendix H.
RQ: Does SynthesizRR address entity diversity? Popularity bias is a phenomenon wherein LLM generations tend to over-represent popular “head” entities. This has been studied for QA tasks (Mallen et al., 2023; Kandpal et al., 2023).
In Figure 4 we see how SynthesizRR eliminates popularity bias across entity types. By sourcing from the long-tail of retrieval results ($k=50$), the generated dataset has much higher entity entropy compared to FewGen. This positions SynthesizRR closer to Gold, which also shows high entity entropy.
@@ 5 Results: Intrinsic Evaluation
SynzthRR 92.0 72.8 87.9 75.2 87.5 89.9
SynzthRR $91.76$ $67.86$ $67.16$ $75.14$ $86.95$ $83.2$
Table 5: MAUVE similarity score ($\uparrow$) using GPT2-XL embeddings. Notation: *32-shot; 3-shot RetrICL; 32-shot Non-RetrICL.
@@
(Dataset size) ($8$K) ($2$K) ($8$K) ($8$K) ($2$K) ($4$K)
Gold - 91.015 78991.01578991.015\,78991.015 789 93.230 76993.23076993.230\,76993.230 769 $82.488$ 81.516 66781.51666781.516\,66781.516 667 $93.06$ $95.2645$ 89.43
Seed - $83.9$ $82.5$ $67.5$ $71.7$ $85.0$ $90.9$ 80.25
Zero-shot
@@
Table 6: Test Accuracy ($\uparrow$) after distilling DeBERTa-v3-Large student from LLaMa-2 Chat 13B and Claude Instant-v1. Contriever was used as the retriever in SynthesizRR. We report the average of 5 runs and rerun in cases where  std. dev. $\geq$6(indicating one or more models failed to converge). The top half considers zero-shot synthesis and bottom half uses in-context learning, and we bold the best result under each paradigm. Notation: *32-shot; 3-shot RetrICL; 32-shot Non-RetrICL.
RQ: How is entity similarity in synthetic data affected by grounding to an in-domain corpus? For the Category task we generate $8$K product reviews and randomly select $8$K Gold examples. In Table 4, we measure entity recall, and find that the occurrence of Gold entities is 100%-140% higher in SynthesizRR than FewGen. The KL divergence of each entity distribution is also lower. We finally consider the entity coverage (unique entities) and entity density (total entities). Compared to Gold, FewGen tends to produce fewer unique entities (places, events, languages, currencies, etc). Each FewGen example also has a lower density of entities, as visible in Table 3. SynthesizRR coverage and density more closely match Gold.
RQ: How distributionally similar are our generated examples and human-written examples? We see from MAUVE scores in Table 5 that zero-shot generations are quite dissimilar in both approaches compared to few-shot methods. Surprisingly, SynthesizRR generations are much more similar to human text than FewGen, despite the fact that nothing in our content sourcing strategy explicitly guides SynthesizRR generations to match the distribution of Gold.
We thus manually inspect generations and discover an interesting pattern which can be attributed to content sourcing. As shown earlier, and in Table 3, the density of entities is higher under SynthesizRR. FewGen produces generations which obey the prompt, but are very bland and do not include specifics. On the other hand, by obtaining information-rich documents, SynthesizRR is able to ground the task inversion step in details of the retrieved article/product. We hypothesise that this improves the MAUVE score towards Gold, which is similarly grounded in specifics.
@@ 6 Results: Student distillation
Table 6 shows the results of training a DeBERTa-v3-Large student on datasets generated by SynthesizRR and FewGen, as well as baselines of tuning on the Gold set and Seed set. In the zero-shot setting, we find that SynthesizRR performs much better than FewGen, despite using the same frozen teacher LLM. Note that SynthesizRR uses in-context examples for retrieval here whereas FewGen does not; our method has some additional supervision here. However, in this setting, we see clear gains during the task inversion stage ($\uparrow$12for LLaMa and $\uparrow$17.6for Claude). Thus, having access to retrieval yields a better final dataset, almost on par with 32-shot FewGen.
@@
With ICL, 3-shot SynthesizRR using the RetrICL strategy trains better students than 32-shot FewGen ($\uparrow$1.3for LLaMa and $\uparrow$3.2for Claude) and Non-RetrICL. We conclude that naively adding ICL examples is not an effective use of the LLM’s context window. Instead, a better content sourcing strategy improves the student distillation, leading to better test performance.
Method Retriever Teacher Self-BLEU-5 ()($\downarrow$)( ↓ ) Entity Entropy ()($\uparrow$)( ↑ ) Mauve ()($\uparrow$)( ↑ ) Accuracy ()($\uparrow$)( ↑ )
(Dataset) LLM AG. IMDb SST-2 AG. IMDb SST-2 AG. IMDb SST-2 AG. IMDb SST-2
Gold - - 17.1 27.9 35.5 6.6 7.5 3.2 - - - 90.8 91.3 88.2
SunGen - GPT2-XL
[…]
15.4
[…]
4.9
[…]
68.7
[…]
84.9
[…]
ReGen BERT - 56.5
[…]
8.1
[…]
68.1
[…]
82.7
[…]
S3 - GPT3.5
$\otimes$
62.2
$\otimes$
$\otimes$
5.7
$\otimes$
$\otimes$
62.0
$\otimes$
$\otimes$
87.1
$\otimes$
AttPmt - GPT3.5-T 39.8
[…]
71.5 6.0
[…]
3.4 52.8
[…]
50.0 79.8
[…]
80.8
Zero-shot
SynzthRR Contr. LLaMa2 29.3 66.3 41.9 7.1 5.7 4.5 89.5 58.5 50.0 85.3 82.9 80.2
SynzthRR Contr. ClaudeV1 31.5 51.5 45.3 6.6 5.3 4.8 94.2 55.9 50.0 85.6 83.6 82.5
SynzthRR BM25 LLaMa2 28.7 62.2 36.5 7.0 5.6 5.1 90.3 60.5 50.0 84.3 74.1 84.4
SynzthRR BM25 ClaudeV1 30.9 50.4 36.9 6.5 5.1 5.4 90.8 53.2 50.0 84.2 79.1 82.6
3-shot RetrICL
SynzthRR Contr. LLaMa2 34.2 62.9 26.3 7.2 5.7 3.8 92.6 72.6 50.0 84.6 84.8 83.8
SynzthRR Contr. ClaudeV1 23.7 38.0 24.6 6.7 5.9 4.3 95.8 58.0 50.0 86.0 86.3 80.6
SynzthRR BM25 LLaMa2 32.0 59.7 25.3 7.2 5.6 4.8 92.5 78.7 50.0 84.3 84.7 84.4
SynzthRR BM25 ClaudeV1 24.6 41.9 26.8 6.7 5.4 4.9 96.0 58.5 50.0 84.1 81.6 82.3
Table 7: Evaluations of synthetic datasets released by prior work. We subsample all to 6K examples (uniformly distributed across classes) before computing metrics as described in §4. Tasks not evaluated by previous authors are denoted by
$\otimes$
while those evaluated without dataset release are marked
[…]
. GPT3.5 is text-davinci-003 whereas GPT3.5-T is gpt-3.5-turbo (OpenAI, 2022), LLaMa2 is 13B Chat version (Touvron et al., 2023a), ClaudeV1 is Instant-V1.2 version (Anthropic, 2023). Accuracy is measured on a DistilBERT student, where we train 5 student models and report the mean accuracy (std. dev. was $\leq 2.0$ in all cases). Within each dataset, we bold the best result.
7 Comparison to previous work
We benchmark SynthesizRR against four prior synthesis methods: (1) SunGen Gao et al. (2023) uses ZeroGen to create 200k synthetic rows and employs a custom bi-level optimization algorithm to weight each instance; (2) ReGen Yu et al. (2023b) utilizes two BERT models, one for retrieval and one as a classifier, to multi-round filter noisy data; (3) S3 Wang et al. (2023a) builds and iteratively enhances a seed dataset by identifying and synthesizing corrections using an LLM; (4) AttrPrompt Yu et al. (2023a) improves dataset diversity and unbiasedness by prompting GPT3.5-Turbo with varied attributes (derived from a human-in-the-loop analysis of each task). Standard zero-shot and few-shot generation baselines were compared in Table 6, so we do not include them here. ZeroGen (Ye et al., 2022a) is similarly excluded.
We benchmark three popular tasks: IMDb Maas et al. (2011), SST-2 Socher et al. (2013) and AG News (Zhang et al., 2015). Previous studies have generated larger datasets ranging from 20k to 200k examples with varying student model hyperparameters, but often lack reports on intrinsic dataset quality, making a fair comparison challenging. Therefore, we independently reproduce these results using the synthetic datasets released by the original authors [^2] ProGen (Ye et al., 2022b) was excluded as it does not release datasets.. Following Yu et al. (2023a), we subsample these datasets to 6k rows, keeping a uniform distribution across classes, and generate the same number of synthetic covariates using SynthesizRR RetrICL (Algorithm 1). For the content sourcing stage of SynthesizRR, we retrieve documents from the CMU Movie Summary corpus Bamman et al. (2013) and RealNews/Dom(Appendix I). We measure accuracy on a DistilBERT student (Sanh et al., 2019; Yu et al., 2023a; Ye et al., 2022a; Gao et al., 2023; Wang et al., 2023a; Ye et al., 2022b), fixing hyperparams to Yu et al. (2023a).
RQ: How does SynthesizRR perform against prior methods on student model accuracy?
Methods like SunGen rely on relatively weak LLM teachers like GPT2-XL Radford et al. (2019) can perform well on topic and sentiment tasks like IMDb, but require a very high data cost (15-30x more synthetic data than SynthesizRR). In Table 7, we observe that when scaled down to 6k rows, the performance deteriorates significantly. We hypothesize that adding the student model into the synthesis process impacts the final classification accuracy, as the dataset becomes specialized to the particular choice of student and does not generalize to other students.
Approaches which use strong instruction-following LLMs like AttrPrompt, S3, and SynthesizRR can achieve similar or better performance with much smaller datasets, as they create high-quality datasets. Prompting techniques like Chain-of-Thought (Wei et al., 2022) used by S3 further improve the task-inversion step (while necessitating higher API costs due to longer output lengths). Chain-of-Thought prompting thus seems like a promising approach to augment SynthesizRR’s task-inversion step.
Table 7 compares diversity (Entity-Entropy, Self-BLEU), and similarity to Gold texts (MAUVE). Only AttrPrompt (Yu et al., 2023a, Appendix E) attempts to improve diversity of the generated text, by templatizing the task inversion instruction with attributes such as style, topic, length:min-words and more. ReGen is the only prior approach to use content sourcing (but not task inversion). These are thus the most relevant baselines for SynthesizRR.
Both ReGen and SynthesizRR achieve very high entity entropy compared to AttrPrompt, underscoring the importance of a content sourcing step. Unlike SynthesizRR, ReGen uses only retrieval without task-inversion, and thus suffers in terms of lexical diversity, MAUVE and student accuracy.
On the other hand, CoT-style prompting (S3) suffers a lack of lexical diversity and similarity to Gold texts, despite strong distillation performance. This is reproduced in AttrPrompt and previously in FewGen, lending evidence to our claim that synthesis without content sourcing tends to produce datasets with lower diversity, which cannot be overcome by complex prompting strategies alone.
Finally, SunGen exhibits high diversity on IMDb, a task for generating sentiment-based movie reviews. Unlike traditional zero-shot generation, SunGen begins by creating a movie with the prompt Movie: followed by generating an example using prompt The movie review in positive sentiment for movie "<Movie>" is: (details in Ye et al. (2022a, Section 4.6)). We posit that this generated movie fulfils a similar purpose to a retrieved context, enhancing the diversity.
8 Related Work
Dataset synthesis using LLMs. Using LLMs to perform task inversion for dataset synthesis has been studied previously. Most use GPT-2XL without fine-tuning (Ye et al., 2022b, a; Gao et al., 2023; Meng et al., 2022; Schick and Schütze, 2021; Jung et al., 2023). Recent work has considered large teacher LLMs such as GPT-3 (West et al., 2022; Honovich et al., 2023; Wang et al., 2023b), PaLM-540B (Hsieh et al., 2023) and chat-tuned LLMs such as gpt-3.5-turbo (Yu et al., 2023a; Yehudai et al., 2024b; Wang et al., 2023a).
For the generation of text classification datasets, class-conditioned prompting is key. Prior approaches investigated zero-shot (Ye et al., 2022a) and iterative few-shot prompting (Ye et al., 2022b), or synthesis using seq2seq LLMs fine-tuned on a curated dataset (Lee et al., 2021). Recently, AttrPrompt (Yu et al., 2023a) established that varying prompt attributes improves diversity. Our work explores adding retrieval contexts as the source of diversity.
Retrieval-augmented generation. Our approach has many of the characteristics of in-context retrieval-augmented generation (RAG) (Lewis et al., 2020; Ram et al., 2023; Huang et al., 2023; Izacard et al., 2023). Previous studies show how RAG bypasses numerous problems associated with generating solely from parametric memory, i.e., heightened bias towards “head” entities (Mallen et al., 2023), lower lexical diversity (Holtzman et al., 2019; Jentzsch and Kersting, 2023), and hallucinated information (Zhang et al., 2023).
Using retrieval-augmented generation for synthesis of classification tasks has not been explored at the instance level. ReGen (Yu et al., 2023b) studies the retrieval-only setting for creation of topic and sentiment datasets, which are simpler than the tasks in our work. Viswanathan et al. (2023) and Gandhi et al. (2024) perform dataset-level retrieval and not instance-level retrieval.
9 Conclusion
Limitations
Appendix A Risks
Although the main goal of our work is to improve text classification, our use of LLMs to generate examples does carry some conceptual risks. By generating news articles to train classifiers on, we run the risk of generating fake news and other harmful content. However, we believe this risk is mitigated by the fact that the final outcome of our system is a classifier: classification models have relatively constrained failure modes (misclassification) compared to text generation models that can mislead users. Furthermore, we do not believe our approach uniquely advances the generation of content like fake news; our advances are largely orthogonal to the technology that brings such risks.
Appendix B Incorporating feedback from distilled student models